OpenStudy (ttop0816):

am i correct?? 1/cosθ+ tan^2θ/cosθ 1 1 + sinθ cos^3θ << sec^3θ

OpenStudy (anonymous):

no u r wrong

OpenStudy (ttop0816):

how come? then would it be 1+sinθ?

OpenStudy (anonymous):

sec ^3(θ)

OpenStudy (anonymous):

ok so in numerator we have 1+ tan^2(θ) = sec ^2(θ) and in denominator we have cos (θ) => 1/ cos (θ) = sec (θ) sec ^2(θ) * sec (θ) = sec ^3(θ)

OpenStudy (ttop0816):

ohhhhhh! i i got the numerator part but i didnt think of 1/cos(θ)= sec(θ) !! thanks a lot (:

OpenStudy (anonymous):

Welcome :D